OpenStudy (anonymous):

solve for x, x^4 - 1 / x^3 = 0

OpenStudy (anonymous):

x^4 - 1/x^3 = 0 multiply both sides by x^3 to get: x^4 - 1=0 x^4=1 x=1

OpenStudy (akashdeepdeb):

x = +-1 Because x^3 goes to the other side and thus x^4 becomes 1 Thus x=+-1 [TWO ANSWERS]